OpenStudy (anonymous):

Can the mean value theorem for integrals be applied to an integral with an average value of 0?

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

The requirement is a continuous function.

OpenStudy (anonymous):

Okay, perhaps I did the problem wrong, but I followed the formula and got two number not in between the interval.

OpenStudy (anonymous):

You'd have something like this:\[ f(c) = \frac{1}{b-a}\int_a^bf(x)~dx\to 0=\frac{1}{b-a}\cdot 0 \]

OpenStudy (anonymous):

Okay, what is the problem?

OpenStudy (anonymous):

f(x) = x3 − 9x on the interval [−1, 1]

OpenStudy (anonymous):

x^3-9x from (-1,1)

OpenStudy (perl):

$$\int_{-1}^{1} x^3-9x = f(c)(1-(-1)) \\ 0 = f(c) * 2 \\ 0= f(c)\\ 0 = c^3 - 9c\\ 0 = c(c^2 - 9)\\ c = 0 \text{ or} ~ c = 3, -3\\ \text {but 3 and -3 are not in the interval (-1,1)}\\ c = 0 \\ \text{0 is in the interval (-1,1)} $$

OpenStudy (anonymous):

Yeah, so in this case, \(f(0) = 0\), meaning \(c=0\).

OpenStudy (perl):

Note that for any odd function f(-x) = -f(x) $$ \int_{-a}^{a} f(x) =0$$

OpenStudy (anonymous):

Okay, I see. Thank you very much.